# Rational Metareasoning for Large Language     Models

### v2

@@ 4.2 Baselines
We illustrate the advantages of our model by comparing its performance to two types of prompting strategies: Direct prompting, where the model is required to provide an immediate answer, and Chain of Thought prompting (CoT), where the model is encouraged to reason through the problem step-by-step before arriving at a solution. Since we are using pretrained models which are not specifically trained for instruction following, we provide five few-shot examples for each task from the unused portion of the training dataset. These examples are carefully chosen to ensure that the length of the reasoning chain matches the perceived difficulty of the question. In addition to these prompting methods, we adopt a finetuning baseline, comparing our method to STaR (Self-Taught Reasoner; Zelikman et al. 2022), which also uses the expert iteration algorithm. While CoT prompting may not yield optimal trajectories, more advanced methods (Yao et al., 2023a; Zheng et al., 2024; Madaan et al., 2023) often increase sequence lengths. Focusing on efficiency, we find CoT’s simplicity ideal for testing our method. Reducing reasoning tokens without performance loss here suggests similar gains in more complex approaches.
@@ 4.3 Training Details
For our experiments, we use Microsoft Phi-2 (Javaheripi et al., 2023), Meta Llama-3-8B (Dubey et al., 2024), and Mistral-7B-v0.3 (Jiang et al., 2023) as the pretrained base models. We have chosen $\gamma=0.1$ to align the distributions of costs and rewards more closely, although we have found the method to be robust to small variations in the choice of this hyperparameter. We sample $K=4$ reasoning chains for each question, using a temperature $t$ of 0.5 and a $top_{p}$ value of 0.9. These parameters are chosen to balance the exploration and exploitation trade-off, allowing us to generate diverse yet relevant reasoning chains. For each iteration $n$, we sample a dataset $\mathcal{D}_{n}$ of size 512 from the union of four training datasets $\mathcal{D}$ described in 4.1. In the self-supervised fine-tuning step, we use a batch size of 16 and a learning rate of 1e-6. We believe that further improvements are possible through a more comprehensive hyperparameter search; however, due to computational constraints, we leave this for future work. Finally, we evaluate all models using greedy decoding to ensure consistent and deterministic output generation. We use pattern matching techniques to extract the answers; an exact match between the generated answer and the ground truth is considered correct.
@@ 5.1 Performance vs Cost
We first evaluate our approach against baselines (Sec. 4.2) across several datasets (Sec. 4.1). Our key criteria are performance (measured by accuracy) and cost (measured by the number of input and output tokens). Our experiments confirm that across all three models and four datasets, our training approach reduces cost while matching or improving performance (see Table 1 for results averaged across datasets; Tables 4 and 5 for per-dataset results; and Appendix E for example reasoning chains). Fig. 1 shows these results for all three models: the y axis compares cost across different baselines and datasets, while the x axis compares performance. We first consider the performance of our baselines: CoT Few-Shot prompting uses a large number of input and output tokens, but yields reasonable performance. In contrast, Direct Few-Shot prompting uses fewer input (and far fewer output) tokens, but yields poor performance on reasoning-intensive datasets (GSM8K and Proofwriter, see Table 4). STaR improves on all of these approaches, using significantly fewer tokens, while achieving comparable or superior performance. Finally, Metareasoning further improves the cost-performance tradeoff by matching STaR’s accuracy while generating 20-37% fewer tokens on average.
@@
phi-2 Direct Few-Shot 50.7 ($\pm$ 1.3) 441.6 ($\pm$ 0.9) 0.0 ($\pm$ 0.0)
CoT Few-Shot 56.6 ($\pm$ 1.3) 1047.3 ($\pm$ 0.9) 193.9 ($\pm$ 4.3)
STaR 63.2 ($\pm$ 1.5) 75.1 ($\pm$ 0.9) 156.9 ($\pm$ 2.4)
Metareasoning 64.7 ($\pm$ 1.4) 75.1 ($\pm$ 0.9) 123.1 ($\pm$ 2.1)
Meta-Llama-3-8B Direct Few-Shot 55.5 ($\pm$ 1.3) 421.3 ($\pm$ 0.9) 0.0 ($\pm$ 0.0)
CoT Few-Shot 63.7 ($\pm$ 1.4) 995.0 ($\pm$ 0.9) 148.4 ($\pm$ 3.1)
STaR 63.9 ($\pm$ 1.3) 73.3 ($\pm$ 0.8) 119.7 ($\pm$ 2.0)
Metareasoning 64.4 ($\pm$ 1.4) 73.3 ($\pm$ 0.9) 95.6 ($\pm$ 2.2)
Mistral-7B-v0.3 Direct Few-Shot 52.7 ($\pm$ 1.4) 498.2 ($\pm$ 1.0) 0.0 ($\pm$ 0.0)
CoT Few-Shot 57.5 ($\pm$ 1.5) 1180.9 ($\pm$ 1.1) 182.7 ($\pm$ 4.3)
STaR 60.1 ($\pm$ 1.3) 83.2 ($\pm$ 0.9) 157.4 ($\pm$ 2.0)
Metareasoning 60.5 ($\pm$ 1.5) 83.2 ($\pm$ 1.0) 114.4 ($\pm$ 2.2)
Table 1: Comparison of different methods based on accuracy and length metrics, averaged across datasets (means with 95% confidence intervals; bold indicates best performing approaches with overlapping 95% intervals; see Tables 4 and 5 for per-dataset results). Metareasoning achieves comparable performance while using significantly fewer input and output tokens compared to STaR or CoT Few-Shot prompting.
Figure 1: Cost and performance. The plot shows accuracy against the logarithm of the weighted sum of input and output tokens (weighted 3:1, as is common in API pricing).
@@ 5.2 Adaptive computation
Section 5.1 demonstrates that our method reduces computational costs on average. But does it actually teach models to reason adaptively (by adjusting reasoning to match task complexity), or just to reason less? To address this question, we first divided our test set $\mathcal{T}$ based on whether or not reasoning was needed to obtain the correct answer. We split the data based on whether Direct Few-Shot on Phi-2 obtained the correct answer. This yielded a split of 4412 “easy” and 2700 “hard” examples. Adaptive methods should use less computation to solve the easy problems. We can empirically compare the results across methods for these two data splits. As shown in Table 2, all models and methods are able to differentiate between hard and easy problems, generating fewer tokens on easier problems. [^1] We note that we specifically chose CoT few shot examples that demonstrated adaptive reasoning length; we expect the effects here to be dependent on the chosen examples. Intriguingly, STaR decreases the difference in reasoning length between hard and easy problems, suggesting that it trains out this adaptive tendency, biasing the model towards using reasoning on all problems. In contrast, Rational Metareasoning increases the difference in reasoning between hard and easy problems, achieving a length reduction of up to 74.4% on Mistral 7B. This indicates that our approach trains models to reason adaptively, helping them recognize when detailed reasoning is necessary and when a shorter response is sufficient.
@@
Phi-2 CoT Few-Shot 145.0 ($\pm$ 4.0) 86.0 ($\pm$ 2.0) 40.7
STaR 135.0 ($\pm$ 4.0) 91.0 ($\pm$ 1.0) 32.6
Metareasoning 122.0 ($\pm$ 4.0) 57.0 ($\pm$ 3.0) 53.3
Meta-Llama-3-8B CoT Few-Shot 132.0 ($\pm$ 3.0) 57.0 ($\pm$ 1.0) 56.8
STaR 126.0 ($\pm$ 3.0) 64.0 ($\pm$ 2.0) 49.2
Metareasoning 99.5 ($\pm$ 2.5) 27.0 ($\pm$ 1.0) 72.9
Mistral-7B-v0.3 CoT Few-Shot 144.0 ($\pm$ 4.0) 77.0 ($\pm$ 3.0) 46.5
STaR 149.0 ($\pm$ 6.0) 98.0 ($\pm$ 1.0) 34.2
Metareasoning 125.0 ($\pm$ 3.0) 32.0 ($\pm$ 1.0) 74.4
Table 2: The length of generated reasoning chains across models and methods (mean number of tokens with 95% confidence intervals). Adaptive methods should maximize the difference in length between the two distributions. Our method (Metareasoning) reduces the overall length and increases the difference in the length distribution between the Hard and Easy splits (as seen in column Length Reduction), demonstrating an improvement in the model’s ability to adapt the reasoning length to the complexity of the task.
5.3 Generalization
Model Method Accuracy (%) $\uparrow$ # Output Tokens $\downarrow$
Phi-2 STaR 47.9 ($\pm$ 1.3) 110.5 ($\pm$ 1.5)
Metareasoning 48.1 ($\pm$ 1.3) 78.7 ($\pm$ 1.5)
Meta-Llama-3-8B STaR 56.3 ($\pm$ 1.1) 75.4 ($\pm$ 1.4)
Metareasoning 54.5 ($\pm$ 1.4) 48.8 ($\pm$ 2.1)
Mistral-7B-v0.3 STaR 54.5 ($\pm$ 1.3) 112.0 ($\pm$ 1.6)
Metareasoning 54.0 ($\pm$ 1.4) 50.6 ($\pm$ 1.5)
Table 3: Comparison of different methods based on accuracy and length metrics in an out-of-distribution setting on the MMLU benchmark. We report the mean with 95% confidence intervals. Metareasoning achieves comparable performance while generating significantly fewer output tokens compared to STaR.
We assess the out-of-distribution generalization of our method using the MMLU benchmark (Hendrycks et al., 2021, see Section 4.1). As presented in Table 3, our approach achieves comparable performance while generating 28.8% to 54.8% fewer tokens compared to STaR. Among the 57 subcategories of MMLU, Fig. 2 highlights four—high school chemistry, mathematics, U.S. history, and macroeconomics—that effectively showcase the strengths of our model. We specifically chose two subjects (U.S. history and macroeconomics) where the use of chain-of-thought (CoT) reasoning appears to be counterproductive compared to Direct Few-Shot prompting, as well as two subjects (mathematics and chemistry) where the benefits of CoT are more pronounced. Consistent with our findings from the four in-domain datasets, we observe that the reduction in output length achieved by our method is inversely related to the benefits gained from intermediate reasoning chains. In the cases of U.S. history and macroeconomics, our method results in a substantial reduction in generated tokens. In contrast, in mathematics and chemistry, the reduction in token count is smaller.
Figure 2: Cost and Performance for Phi-2 Across Selected MMLU Domains. Left panel: Input and output token distributions for different methods and domains. Right panel: While CoT improves performance in subjects like mathematics and chemistry, it appears less effective in history and macroeconomics. The latter domains show a greater reduction in output tokens (35% for history and 37% for macroeconomics), compared to the smaller reductions seen in chemistry (20%) and mathematics (12%), where intermediate reasoning chains provide more substantial benefits.
@@ 6.2 Reasoning in LLMs
Techniques such as Chain of Thought (CoT) and related methodologies (Wei et al., 2023; Yao et al., 2023a; Yasunaga et al., 2024; Madaan et al., 2023; Zheng et al., 2024) have proven effective at enhancing LLM performance across a wide range of tasks. CoT boosts LLMs’ performance on complex reasoning by guiding them through a series of intermediate reasoning steps, increasing inference costs to improve task performance. This method can be implemented through in-context learning (Wei et al., 2023), prompting (Kojima et al., 2023), or training (Li et al., 2023). The benefits of CoT can be attributed to both a greater computation depth (Goyal et al., 2024; Pfau et al., 2024) and the semantic values of the thought tokens, which function as intermediate variables in the computation of the answer (Prystawski et al., 2023). However, recent studies have raised concerns regarding the meaningfulness of such reasoning chains in reaching the target solution, and whether models effectively utilize them to solve tasks (Turpin et al., 2023; Paul et al., 2024; Sprague et al., 2024). We further demonstrate that standard prompting and training methods fail to teach the model to use CoT purposefully, resulting in inefficient inference. Reasoning can also be used to bootstrap language models. Self-improving techniques (Huang et al., 2022; Zelikman et al., 2022) consist of generating reasoning chain-augmented answers for unlabeled questions and fine-tuning the LLM using those self-generated solutions as target outputs. Similar techniques, such as ReST (Gulcehre et al., 2023), can also be used to better align LLMs with human preferences and needs. Our approach builds on these techniques to optimize the inference cost of reasoning chains in addition to their task performance. Another related work, Quiet-STaR (Zelikman et al., 2024), has improved LLM reasoning performance by modifying pretraining to generate intermediate s between tokens. While effective for downstream tasks, it increases computational costs by generating reasoning chains at every step, even when unnecessary. More recently, chat models that “think before answering” have been developed (OpenAI, 2024), using inference-time computation to enhance their outputs. Although these models outperform others, they expend more computational resources on all tasks, even when it may not be necessary. Our method could be incorporated into their training process to help the model determine when this additional computation is genuinely beneficial.

### v3

@@ 4.2 Baselines
We illustrate the advantages of our model by comparing its performance to two types of prompting strategies: Direct prompting, where the model is required to provide an immediate answer, and Chain of Thought prompting (CoT), where the model is encouraged to reason through the problem step-by-step before arriving at a solution. Since we are using pretrained models which are not specifically trained for instruction following, we provide five few-shot examples for each task from the unused portion of the training dataset. These are the same examples that are used to guide the model training during the first iteration. In addition to these prompting methods, we adopt as a finetuning baseline the most common reasoning bootstrapping method, STaR (Self-Taught Reasoner; Zelikman et al. 2022), which also uses a variation of Expert Iteration. In particular, this method was the first to adopt outcome-based rewards to enable models to self-improve: given a set of problems and solutions, it trains the model iteratively on reasoning chains that lead to the correct solution.
As an upper bound, we also tested the instruction-tuned version of the models, which have been fine-tuned on a massive amount of annotated chain-of-thought data, followed by RLHF (Dubey et al., 2024). Finally, while CoT prompting may not yield optimal trajectories, more advanced methods (Yao et al., 2023a; Zheng et al., 2024; Madaan et al., 2023) often increase sequence lengths substantially, and are therefore less efficient. We focus on CoT for its simplicity – reducing reasoning tokens while gaining performance for CoT suggests similar gains for more complex approaches.
@@ 4.3 Training Details
For our experiments, we use Meta Llama-3.2-3B and Llama-3.1-8B (Dubey et al., 2024) as the pretrained base models. We have chosen $\gamma=0.1$ to align the distributions of costs and rewards more closely, although we have found our method to be robust to small variations in the choice of this hyperparameter. We sample $K=4$ reasoning chains for each question, using a temperature $t$ of 0.5 and a $top_{p}$ value of 0.9. These parameters are chosen to balance exploration and exploitation, allowing us to generate diverse yet relevant reasoning chains.
For greater efficiency during training, instead of using the entire dataset from the start, we begin by sampling a dataset $\mathcal{D}_{n}$ of size $T=1024$ from the union of the four training datasets $\mathcal{D}$ described in 4.1. We then progressively increase the dataset size by $T$ with each subsequent iteration (until it reaches the size of $\mathcal{D}$) allowing the model to encounter new examples gradually. In the self-supervised fine-tuning step, we use a batch size of 16 and a learning rate of 1e-5. We execute five iterations of the training algorithm. We believe that further improvements are possible through a more comprehensive hyperparameter search; however, due to computational constraints, we leave this for future work. Finally, we evaluate all models using greedy decoding to ensure consistent and deterministic output generation. We use pattern matching techniques to extract the answers; an exact match between the generated answer and the ground truth is considered correct.
@@ 5.1 Performance vs Cost
We first evaluate our approach against baselines (Sec. 4.2) across several datasets (Sec. 4.1). Our key criteria are performance (measured by rescaled accuracy) and cost (measured by the number of input and output tokens). Our experiments confirm that across all models and datasets, our training approach reduces cost while matching or improving performance (see Table 1 for results averaged across datasets; Tables 4 and 5 for per-dataset results; and Appendix C for example reasoning chains). Fig. 2 shows these results for all models.
Figure 2: Cost and performance. Accuracy is plotted against the output tokens. Our method, RaM, eliminates the need for few-shot prompting (reducing input tokens) and trains the model to use fewer reasoning tokens than STaR.
@@
Llama-3.2-3B Direct Few-Shot 33.5 ($\pm$ 1.4) 291.7 ($\pm$ 0.9) 0.0 ($\pm$ 0.0)
CoT Few-Shot 34.8 ($\pm$ 1.8) 1080.5 ($\pm$ 0.9) 261.9 ($\pm$ 4.8)
STaR 47.4 ($\pm$ 2.0) 73.2 ($\pm$ 0.9) 222.7 ($\pm$ 3.5)
RaM 50.1 ($\pm$ 1.8) 73.2 ($\pm$ 0.8) 170.8 ($\pm$ 2.9)
Llama-3.2-3B-Instruct CoT Prompting 53.2 ($\pm$ 1.8) 157.2 ($\pm$ 0.9) 259.1 ($\pm$ 3.4)
Llama-3.1-8B Direct Few-Shot 45.8 ($\pm$ 1.5) 291.7 ($\pm$ 0.9) 0.0 ($\pm$ 0.0)
CoT Few-Shot 56.4 ($\pm$ 1.9) 1080.5 ($\pm$ 0.9) 253.6 ($\pm$ 4.2)
STaR 61.9 ($\pm$ 1.6) 73.2 ($\pm$ 0.9) 215.1 ($\pm$ 3.3)
RaM 63.5 ($\pm$ 1.7) 73.2 ($\pm$ 0.9) 146.2 ($\pm$ 3.3)
Llama-3.1-8B-Instruct CoT Prompting 60.9 ($\pm$ 1.7) 157.2 ($\pm$ 0.9) 269.9 ($\pm$ 3.8)
Table 1: Comparison of different methods based on rescaled accuracy and length metrics, averaged across datasets (means with 95% confidence intervals; bold indicates best performing approaches with overlapping 95% intervals). RaM achieves better performance while using significantly fewer input and output tokens compared to STaR or CoT Few-Shot.
We first consider the performance of our baselines: CoT Few-Shot prompting uses a large number of input and output tokens, but yields reasonable performance. Direct Few-Shot prompting uses fewer input (and far fewer output) tokens, but yields poor performance on reasoning-intensive datasets (GSM8K and Proofwriter, see Table 4). STaR improves on these approaches, using significantly fewer tokens while achieving comparable or superior performance.
RaM further improves the cost-performance tradeoff by generating 23-32% fewer tokens on average compared to STaR, generally with higher accuracy. Interestingly, our findings align with the conclusions of a meta-analysis of over 100 papers (Sprague et al., 2024), in finding that explicit chains help mainly on maths or logic tasks and bring marginal gains elsewhere.
Finally, it is worth noting that while the performance of the instruction-tuned model (CoT Instruct) is generally higher than that of the pretrained model, the length of the reasoning chains was generally much higher and not meaningfully adaptive to task complexity.
@@ 5.2 Adaptive computation
Section 5.1 demonstrates that our method reduces computational costs on average. But does it actually teach models to reason adaptively (by adjusting reasoning to match task complexity), or just to reason less? To address this question, we first divided our test set $\mathcal{T}$ based on whether or not reasoning was needed to solve the task. In particular, we split the data based on whether Direct Few-Shot obtained the correct answer (“easy split”) or not (“hard split”). Adaptive methods should use less computation to solve the easy problems. We can empirically compare the results across methods for these two data splits.
As shown in Table 2, all models and methods are able to differentiate between hard and easy problems, generating fewer tokens on easier problems. The instruction-tuned model (CoT Instruct) did not effectively adapt response lengths to align with task complexity in our experiments. While STaR seems to improve this ratio, RaM increases the difference in reasoning between hard and easy problems, achieving a length reduction of up to 50.3% on Llama-3.2-3B. This indicates that RaM trains models to reason adaptively, helping them recognize when detailed reasoning is necessary and when a shorter response is sufficient.
5.3 Generalization
We assess the out-of-distribution generalization of the length reduction of RaM with respect to STaR using the public test set of the MMLU-CF benchmark (Hendrycks et al., 2021, see Section 4.1). As shown in Table 3, RaM achieves similar performance while generating 28% to 36% fewer tokens than STaR.
Looking at subdomains, Fig. 3 shows the ratio between the output length of RaM with respect to STaR. We can see that subjects that require multi-step reasoning (for example, physics or engineering) exhibit smaller reductions because they need detailed intermediate steps. In contrast, tasks focused primarily on recalling factual information tend to show a more pronounced length reduction.
@@ 5.3 Generalization
Llama-3.2-3B CoT Few-Shot 210.0 ($\pm$ 3.0) 170.0 ($\pm$ 3.0) 19.0
STaR 195.0 ($\pm$ 5.0) 128.0 ($\pm$ 1.0) 34.4
RaM 145.0 ($\pm$ 3.0) 72.0 ($\pm$ 2.0) 50.3
Llama-3.2-3B-Instruct CoT Prompting 222.0 ($\pm$ 4.0) 221.0 ($\pm$ 3.0) 0.5
Llama-3.1-8B CoT Few-Shot 197.5 ($\pm$ 3.5) 151.0 ($\pm$ 1.0) 23.5
STaR 202.0 ($\pm$ 5.0) 129.5 ($\pm$ 2.5) 35.9
RaM 139.0 ($\pm$ 3.0) 89.0 ($\pm$ 1.0) 36.0
Llama-3.1-8B-Instruct CoT Prompting 228.0 ($\pm$ 6.0) 226.0 ($\pm$ 2.0) 0.9
Table 2: Length of generated reasoning chains across models and methods (mean number of tokens with 95% confidence intervals). Adaptive methods should maximize the difference in length between the two distributions. RaM reduces the overall length and increases the difference in the length distribution between the Hard and Easy splits (Length Reduction), demonstrating an improvement in the ability to adapt reasoning length to task complexity.
Llama-3.2-3B STaR 27.0 ($\pm$ 2.4) 75.1 ($\pm$ 0.9) 150.6 ($\pm$ 4.1)
RaM 29.5 ($\pm$ 2.5) 75.1 ($\pm$ 0.9) 96.3 ($\pm$ 3.7)
Llama-3.1-8B STaR 40.2 ($\pm$ 2.4) 75.1 ($\pm$ 0.8) 153.6 ($\pm$ 4.0)
RaM 36.4 ($\pm$ 2.4) 75.1 ($\pm$ 0.9) 110.6 ($\pm$ 3.8)
Table 3: Comparison of STaR and RaM based on rescaled accuracy and length metrics in an out-of-distribution setting on the MMLU-CF benchmark. We report the mean with 95% confidence intervals. RaM has similar performance but generates fewer output tokens.
Figure 3: Ratio of output length reduction. RaM ’s output length relative to STaR’s, for Llama-3.2-3B. Tasks requiring heavier reasoning tend to exhibit a lower length reduction ratio, whereas those relying more on knowledge display a more pronounced length reduction.
@@ 6.1 Reducing inference costs
Finally, the concurrent “long-to-short” literature deliberately compresses reasoning paths while preserving accuracy. (Arora & Zanette, 2025) shows that adding a length penalty to verifiable rewards can shorten the reasoning length of Large Reasoning Models in mathematical settings. Here, we focus on bootstrapping the reasoning process, teaching pretrained language models to reason efficiently from scratch by developing a reward that is agnostic to the answer format. CoT-Valve (Ma et al., 2025) and Claude 3.7 (Anthropic, 2025) enable external control over the length of the reasoning chains (and therefore the computation budget), allowing a model to produce both long and short reasoning chains. Unlike these methods, our method enables the model to independently adapt the reasoning length according to the difficulty of the task. Concurrent work (Wu et al., 2025; Luo et al., 2025) also shows that reasoning length compression can also be achieved by merging LLMs and Large Reasoning Models. Similarly, Synergy-of-Thoughts (Shang et al., 2024) employed models of different sizes to adapt to the complexity of the task at hand. Our approach differs from all of the above on two fronts. First, it keeps a single set of parameters: there is no external controller, merged checkpoint, or auxiliary model. Second, adaptivity is learned through a principled Value-of-Computation reward, not by hard length constraints or heuristic escalation.
@@ 6.2 Reasoning in LLMs
Techniques such as Chain of Thought (CoT) and related methodologies (Wei et al., 2023; Yao et al., 2023a; Madaan et al., 2023; Zheng et al., 2024) have proven effective at enhancing LLM performance across a wide range of tasks. CoT boosts LLMs’ performance on complex reasoning by guiding them through a series of intermediate reasoning steps, increasing inference costs to improve task performance. This method can be implemented through in-context learning (Wei et al., 2023), prompting (Kojima et al., 2023), or training (Li et al., 2023). The benefits of CoT can be attributed to both a greater computation depth (Goyal et al., 2024; Pfau et al., 2024) and the semantic values of the thought tokens, which function as intermediate variables in the computation of the answer (Prystawski et al., 2023). However, recent studies have raised concerns regarding the meaningfulness of such reasoning chains in reaching the target solution, and whether models effectively utilize them to solve tasks (Turpin et al., 2023; Paul et al., 2024; Sprague et al., 2024). We further demonstrate that standard prompting and training methods fail to teach the model to use CoT purposefully, resulting in inefficient inference. Moreover, reasoning in situations where it is unnecessary may also harm model performance Liu et al. (2024).
Reasoning can also be used to bootstrap language models. Self-improving techniques (Huang et al., 2022; Zelikman et al., 2022) consist of generating reasoning chain-augmented answers for unlabeled questions and fine-tuning the LLM using those self-generated solutions as target outputs. Similar techniques, such as ReST (Gulcehre et al., 2023), can also be used to better align LLMs with human preferences and needs. Our approach builds on these techniques to optimize the inference cost of reasoning in addition to task performance.
Another related work, Quiet-STaR (Zelikman et al., 2024), has improved LLM reasoning performance by modifying pretraining to generate intermediate reasoning sequences between tokens. While effective for downstream tasks, it increases computational costs by generating reasoning chains at every step, even when unnecessary. More recently, chat models that “think before answering” have been developed (OpenAI, 2024; Anthropic et al., 2025; DeepSeek-AI et al., 2025), using inference-time computation to enhance their outputs. Although these models outperform others, they expend more computational resources even when it may not be necessary. Our method could be incorporated into their training process to help the model determine when this additional computation is genuinely beneficial.
7 Limitations and Future Work
While our experiments show that rational metareasoning can successfully reduce inference costs in large language models, some limitations remain. First, our approach has only been tested on well-established datasets in science, commonsense reasoning, logic, and mathematics. Its effectiveness in other contexts has yet to be demonstrated. One particularly relevant example is the agentic setting, where LLMs act autonomously in complex digital environments (Yao et al., 2023b; Schick et al., 2023). Adapting our method to this context would require incorporating the cost of tool use (e.g., API calls) into the reward function to encourage the model to minimize unnecessary resource consumption. Second, while the VoC reward function is highly flexible and can be readily combined with various learning algorithms, our work focused on Expert Iteration (Havrilla et al., 2024) due to its sample efficiency for reasoning tasks. Extending the VoC reward framework to other training approaches like PPO or DPO remains an open research area. Evaluating whether VoC-based rewards maintain or even improve performance in these settings would further clarify the robustness and generalizability of our approach.